OpenStudy (anonymous):

Can someone show me how to simplify this radical expression? Much appreciated! ^6√64a^3b^3

OpenStudy (anonymous):

^6√64a^3b^3 doesn't make sense. in the beginning ^6 has no value in front of it?

OpenStudy (anonymous):

its an exponent. straight from my book

OpenStudy (anonymous):

there is nothing in front of ^6?

OpenStudy (anonymous):

a radical

OpenStudy (anonymous):

it cant have no value in front of it... it just doesnt make sense. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I think YourRoyalFatness means that the expression is \[\large \sqrt[6]{64a^3b^3}\]

OpenStudy (anonymous):

now it makes sense.

OpenStudy (anonymous):

yes jim you are right!

OpenStudy (anonymous):

i am so stumped on this problem

jimthompson5910 (jim_thompson5910):

that's correct str99, you can go further to say this \[\large \sqrt[6]{64a^3b^3}\] \[\large \sqrt[6]{64}*\sqrt[6]{a^3b^3}\] \[\large 2\sqrt[6]{a^3b^3}\] \[\large 2(a^3b^3)^{1/6}\] \[\large 2(a^{3/6}*b^{3/6})\] \[\large 2(a^{1/2}*b^{1/2})\] \[\large 2(ab)^{1/2}\] \[\large 2\sqrt{ab}\]

OpenStudy (anonymous):

thanks! it matches the answer in the back of the book

zepdrix (zepdrix):

@str99 That notation was maybe a little confusing. The carot wasn't being used to show an exponent. It was showing the degree of the root. :)

OpenStudy (anonymous):

You are right. Thanks for correcting me